OpenStudy (anonymous):

Does anybody know a algebraic formula I could make out of this problem? Currently Dan has $600 in his bank account, while Kerri has $500 in hers. Every Friday afternoon Dan withdraws $15. At the same time Kerri withdraws $10 knowing that she can't spend as much as Dan. After how many weeks will their account balances be equal?

OpenStudy (anonymous):

\[600-15x=500-10x\] where x is the number of weeks.

OpenStudy (anonymous):

And how can I find the number of weeks?

OpenStudy (anonymous):

solve for x. you can add 15 x to both sides to get \[600=500+5x\] subtract 500 to get \[100=5x\] divide by 5 get \[x=20\]

OpenStudy (anonymous):

Thanks for helping a 6th grader. I really appreciate it! I hope I see you again in the future satellite!

OpenStudy (anonymous):

Oh, and also I know it's a stupid question but where did you get 5x?

OpenStudy (anonymous):

i had \[-10x\] on the right and \[-15x\] on the left. so i did this \[600-15x+15x=500-10x+15x\] \[600=500+5x\] because \[-10x+15x=5x\]

OpenStudy (anonymous):

yw